Consider the reaction of NO with ozone (a reaction present in the stratosphere that is responsible for ozone depletion):

NO + O3 -> NO2 + O2

You set up an experiment in which you have two flasks separated by a valve. Each flask has a pressure of 10 millibars (or 7.50 mmHg) and a temperature of -35.0C, mimicking the conditions in the middle of the stratosphere. One flask contains 0.0200 grams of nitrogen oxide. The second flask contains ozone and has a volume of 3.00 L. You open the valve and allow the gases to react, creating nitrogen dioxide and oxygen gas.

After the reaction has gone to completion, what is the total pressure inside your two-flask apparatus? Please provide your answer in mmHg.
1
Expert's answer
2020-12-14T14:35:04-0500

Moles of "NO_2=" "0.0200g\over 46.0g/mol" "=0.00043mol"

Moles of "O_3=" "3000g\over 48g/mol" "=62.5mol"

From the ideal gas law;

"P_{NO_2}=" "nNO_2RT \over V"

"=" "0.00043mol \times {0.08206{atm.L\over mol .K}\times 238K}\over 3L" "=0.00279atm"

"1atm =760mmHg"

"=2.1204mmHg"

"P_{O_3}=" "62.5mol\times {0.08206{atm.L\over mol.L}\times 238K}\over3L" "=406.88atm"

"=309228.88mmHg"

"P_{total}=P_1+P_2"

"=2.1204+309228.88=309,230.92mmHg"


.